A+B+C=12;if A+B=13 then find the value of C.(note C is a normal number)

Grade:12

1 Answers

Neelesh Kumar
6 years ago
A+B+C=12
…..............................
A+B=13
…....................................
C=12-(A+B)
….........................
C=12-13
…...........................
C=-1.............